A recipe calls for 1 1/3 tablespoons of butter and serves 6 people.If you wish
to cook this recip for yourself,how much butter should you use?

Answer by AnlytcPhil(1810)

A recipe calls for 1 1/3 tablespoons
of butter and serves 6 people.If you
wish to cook this recipe for yourself,
how much butter should you use?

You would divide 1 1/3 by 6.

(1 1/3) ÷ 6 =

(4/3) ÷ (6/1)

(4/3) × (1/6)

   4/18

   2/9 tablespoons

Since there are no measuring 
spoons for ninths, we use the
fact that 1 tablespoon = 
3 teaspoons

   2/9 tablespoon =

   (2/9) × (3 teaspoons) =

   (2/9)×(3/1) teaspoon =

    6/9 teaspoon =

    2/3 teaspoon  

There are no third teaspoon
measuring spoons either so you
have to measure a teaspoon and
take a litle bit out.

Edwin